(3x-7)^2=8
can someone show me what to do??
Thanks
solve for x

Answer by stanbon(48568)

(3x-7)^2=8
Take the square root of both sides to get:
3x-7=2sqrt2
3x=(2sqrt2)+7
x=[2sqrt2+7]/3
-------------
Another way to do it is as follows:
(3x-7)^2-(sqrt8)^2=0
Factor to get:
[3x-7-sqrt8)(3x-7+sqrt8)]=0
Then:
3x-7=sqrt8 pr 3x-7=-sqrt8
x=[sqrt8+7]/3 or x=[-sqrt8+7]/3
x=[2sqrt2+7[/3 or x=[-2sqrt2+7]/3
-----------
Comment: Notice that the 1st way of doing the
problem results in only one of the possible
solutions. The 2nd way is the better way
Cheers,
Stan H.